# Monitoring-feedback induced entanglement relaxations in a tilted free fermionic chain

### v1

@@ IV Entanglement relaxations and dynamical transition
Now, we delve into the dynamics of entanglement relaxation by examining the bipartite entanglement entropy, denoted as $\overline{S}_{L/2}$, across various $L$ and $\Delta$. Under OBCs, in FIG. 3 (a), $\overline{S}_{L/2}$ first sharply increases, then after the transient metastable value, slowly declines until the steady value irrespective of the presence or absence of the tilted potential. The entanglement relaxation with the tilted potential ($\Delta=0.6$) requires more time than the case without it ($\Delta=0$). Notably, there is a cross point during the decline of $\overline{S}_{L/2}$ for different $L$ and the same $\Delta$. We emphasize them with dashed lines in zoomed insets, called transition points. The transition point separates the log-law and area-law regions. From the example of FIG. 3 (a) inset, in the left of the transition point the relationship between $\overline{S}_{L/2}$ and $\log{L}$ is a linear increase; but in the right, $\overline{S}_{L/2}$ remains constant regardless of $\log{L}$. For $\Delta=0$, separated time (rescaled time ${\rm time}/L$) is about $0.79$, which is consistent with the results in Ref. [70]; for $\Delta=0.6$, the time is approximately $1.47$. As the duration of the entanglement-relaxation process increases, the transition point is delayed, corresponding to the zone below the no exact edge shown in FIG. 1 (b) which are separated from the area law region with a fuzzy boundary. We have noticed that if $\Delta$ is significant enough and entanglement relaxations are prolonged enough, there might be no dynamical entanglement phase transition. For this case, $\overline{S}_{L/2}$ stays in the area-law region after an initial rapid increase. This novel phenomenon is demonstrated in the zone located above the non-exact edge in FIG. 1 (b).
Figure 4: Under OBCs, the dynamical phase diagram of $\overline{S}_{L/2}$ in the plane $\left(\text{time}/L,\Delta\right)$ for $L=192$ for relatively weak $\Delta$.
@@
In FIG. 3 (c), considering $\Delta=0$ under PBCs, after a brief increase, $\overline{S}_{L/2}$ remains stable, which is log-law for the steady value. From the inset, $\overline{S}_{L/2}$ increases linearly with respect to $\log{L}$, and different-time relationships are basically the same. If $\Delta=0.6$, the behavior of $\overline{S}_{L/2}$, starts with a sharp increase, moves to metastable value, and finally declines to a steady value, which is the same as the case under OBCs, in terms of the entanglement relaxations and entanglement transitions due to the pseudo edge.
@@
The maximum trajectory averaged bipartite entanglement entropy $\overline{S}_{L/2}^{max}$ could be found in metastable value. FIG. 3 (b) and (d) illustrate how $\overline{S}_{L/2}^{max}$ changes with $\Delta$ for different $L$, under OBCs and PBCs separetely. In FIG. 3 (b), $\overline{S}_{L/2}^{max}$ increases with $\Delta$ until the peak, about $0.4$, then decreases. $\overline{S}_{L/2}^{max}$ undergoes a transition from log law to area law, and we provide some illustrative instances of the different linear relationships between the $\overline{S}_{L/2}^{max}$ and $\log{L}$ in the inset of FIG. 3 (b). FIG. 3 (d) shows the PBCs case, we find that only $\Delta=0$ there is a huge difference between PBCs and OBCs because of the absence of the pseudo edge. We can also see the transition from log law to area law as $\Delta$ increases.
We have constructed the dynamical phase diagram depicted in FIG. 4 by studying $\overline{S}_{L/2}$ at various times and across different values of $\Delta$. This figure can be viewed as an enlarged detail of FIG. 1 (b), specifically focusing on the low $\Delta$ zone. In FIG. 3 (a), the entanglement relaxation processes described by the darkest green and orange solid lines can also be found in FIG. 4. Upon observation, it is evident that the duration of entanglement relaxation extends as $\Delta$ increases. Concurrently, the transition point that distinguishes between the log-law and area-law regions remains consistent but shifts to the right. The postponed transition point can be attributed to the fact that the tilted potential decelerates the motion of the particles. When the motion is sufficiently slowed, the outcome is depicted in FIG. 1(b).
Figure 5: $L=64$ and OBC. (a) shows the variation of the trajactory averaged $\overline{v}$ over rescaled time. (b) illustrates that the relaxtion acceration $\overline{RA}$ versus $\Delta$. The data of $\overline{RA}$ are from the linear fit of the decaying $\overline{v}$ in (a).
V The interplay between bulk dynamics and edge effect
@@
From the perspective of particle dynamics, we conduct discussions of entanglement relaxation. Under OBCs, we have chosen the trajectory averaged velocity of particles, represented as $\overline{v}$ [70], as our key observable. In FIG. 5 (a), particles are initially given a velocity directed to the left. This velocity then decays linearly over time, eventually to stable zero, which marks the end of the relaxation process. Subsequently, the particles will no longer move. The decay of $\overline{v}$ is primarily due to the edge effect. Furthermore, as $\Delta$ increases, the initial velocity of the particles decreases, resulting in a longer relaxation time required for the system to return to equilibrium. This indicates that $\Delta$ holds the bulk dynamics of the particles.
We have determined the slope of the decay for $\overline{v}$ for different $L$ and depicted these values in FIG. 5 (b), which represent the averaged relaxation acceleration, denoted as $\overline{RA}$. It is observed that $\overline{RA}$ decreases as $\Delta$ increases. The phenomenon can be explained by the fact that the smaller the initial velocity of the particles, the longer time it takes for them to reach the edge, and the edge effects exert less influence. In other words, the tilted potential governs the bulk dynamics and influences the relaxation time to equilibrium.
VI Conclusion and outlook
In summary, we examine the entanglement relaxation dynamics induced by generalized monitoring in a tilted free fermion chain under OBCs and PBCs. Under small tilts, for trajectory averaged bipartite entanglement entropy $\overline{S}_{L/2}$ in rescale time ${\rm time}/L$, there is an entanglement phase transition separating the log-law (induced by weak measurement) and area-law (induced by feedback-driving skin effect) regions and such behavior will be delayed as the increase of $\Delta$. However, under strong tilts, only the area-law region exists, and there may be no exact transition behavior, regardless of how long ${\rm time}/L$ is. On the other hand, the tilted potential can result in the pseudo edge for PBCs, which makes the trajectory averaged density distribution $\overline{\langle\hat{n}_{l}\rangle}$ predominantly concentrate on the left half of the chain just like corresponding condition under OBCs which should have been uniformly distributed across all sites. As long as there is a certain degree of the tilted potential, $\overline{S}_{L/2}$ performs the same under OBCs and PBCs. Thus, the pseudo edge under PBCs and the edge under OBCs play the same role in the entanglement relaxation. Edge effects influence the motion of particles, causing the velocity of particles decreases uniformly. The tilted potential restricts the initial velocity of particles. The interplay between them can be concluded that the smaller the initial velocity of the particles, the later most particles reach the edge, and the weaker the influence of boundary effects. Moreover, the quantum jump with the edge feedback could restrict the pseudo-edge effect to some extent under PBCs.
These findings not only expand the understanding of how feedback influences measurement-induced entanglement phase transitions but also enhance the knowledge of the non-Hermitian skin effect [85] with quantum jumps. In the future, it will be interesting to both re-examine various types of non-Hermitian skin effects driven by feedback monitoring, such as the critical non-Hermitian skin effect [86, 87] without quantum jumps, higher-order skin effects [88], and occupation-particle-dependent skin effects [89], and to explore a universal understanding of feedback-monitored entanglement relaxation.
@@ Acknowledgements.
We thank the helpful discussions with Ze-Chuan Liu and Yu-Peng Wang. JXZ acknowledges the National Natural Science Foundation of China (Grant No. 11874316 and No.12374046), the National Basic Research Program of China (Grant No. 2015CB921103), and the Program for Changjiang Scholars and Innovative Research Team in University (Grant No. IRT13093).

### v2

@@ III.2 Entanglement relaxations and dynamical phase transitions
Now, we delve into the dynamics of the entanglement relaxation by examining the bipartite entanglement entropy, denoted as $\overline{S}_{L/2}$, across various $L$ and $\Delta$. Under OBCs, in FIG. 2 (e), $\overline{S}_{L/2}$ first sharply increases, then after the transient metastable value, slowly declines until the steady value irrespective of the presence or absence of the Wannier-Stark ladder. The entanglement relaxation with the Wannier-Stark ladder ($\Delta=0.6$) requires more time than the case without it ($\Delta=0$). Notably, there is a cross point during the decline of $\overline{S}_{L/2}$ for different $L$ and the same $\Delta$. We emphasize them with dashed lines in zoomed insets, called transition points. The transition point separates the log-law phase from the area-law phase. The inset of FIG. 2 (e) presents several examples, illustrating the relationship between $\overline{S}_{L/2}$ and $\log{L}$ when $\Delta$ is 0.0 and 0.6, respectively. In the left of the transition point the relationship between $\overline{S}_{L/2}$ and $\log{L}$ is a linear increase; but in the right, $\overline{S}_{L/2}$ remains constant regardless of $\log{L}$. For $\Delta=0$, the critical time (rescaled time $t_{c}/\tau$) is about $0.79$, which is consistent with the results in Ref. [1]; for $\Delta=0.6$, $t_{c}/\tau$approx $1.45$. As the duration of the entanglement-relaxation process increases, the transition point is delayed, corresponding to the lower zone of FIG. 1 (b) which comprises both the log-law and area-law regions. We have noticed that if $\Delta$ is significant enough and the evolution is slow enough, we suggest that there might not exist the transition point based on long-term observations. We provide examples in the Appendix C to demonstrate this viewpoint. For this case, $\overline{S}_{L/2}$ stays in the area-law phase after an initial rapid increase. This novel phenomenon is demonstrated in the upper zone of FIG. 1 (b).
@@
In FIG. 2 (h), considering $\Delta=0$ under PBCs, after a brief increase, $\overline{S}_{L/2}$ remains stable, which is log-law for the steady value. From the green solid line of FIG. 2 (h), $\overline{S}_{L/2}$ increases linearly with respect to $\log{L}$, and different-time relationships are basically the same.
If $\Delta=0.6$, the behavior of $\overline{S}_{L/2}$ starts with a sharp increase, moves to metastable value, and finally declines to a steady value, which is the same as the case under OBCs, in terms of the entanglement relaxations and entanglement transitions due to the pseudo edge.
Figure 3: The evolution of different observables under various $\Delta$ values. The columns of the panels correspond to $\Delta=0.0$, $0.2$, $0.4$ and $0.6$, respectively. Panels (a)-(d) display the trajectory-averaged mutual information $\overline{I}_{AB}$. The selected subsystems $A$ and $B$, each with a size of $L/8$ and separated by $L/2$, are depicted in the schematic of panel (a). Panels (e)-(h) describe the evolution of the average probability $\overline{f}_{\text{skin}}$ of the evolved state being in the ideal many-body state over time. Panels (i)-(h) are about $\overline{\|D\|}_{F}$, describing the evolution of the correlations between different subsystems under the system bipartition. The curves in the figure range from light to dark, corresponding to system sizes from small to large.
@@
The maximum trajectory-averaged bipartite entanglement entropy $\overline{S}_{L/2}^{\text{max}}$ could be found in metastable value. FIG. 2 (f) and (i) illustrate how $\overline{S}_{L/2}^{\text{max}}$ changes with $\Delta$ for different $L$, under OBCs and PBCs respectively. In FIG. 2 (f), $\overline{S}_{L/2}^{\text{max}}$ increases with $\Delta$ until the peak, about $0.4$, then decreases. $\overline{S}_{L/2}^{\text{max}}$ undergoes a transition from log-law to area-law, and we provide some illustrative instances of the different linear relationships between the $\overline{S}_{L/2}^{\text{max}}$ and $\log{L}$ in the FIG. 2 (f). FIG. 2 (i) shows the PBCs case, we find that only $\Delta=0$ there is a huge difference between PBCs and OBCs because of the absence of the pseudo edge. We can also see the transition from log-law phase to area-law phase as $\Delta$ increases.
@@
From the perspective of particle dynamics, we conduct discussions of the entanglement relaxation. Under OBCs, we have chosen the trajectory-averaged velocity of particles, represented as $\overline{v}$ [1], as our key observable. In FIG. 2 (g), particles are initially given a velocity directed to the left. This velocity then decays linearly over time, eventually to stable zero, which marks the end of the relaxation process. Subsequently, the particles will no longer move. The decay of $\overline{v}$ is primarily due to the edge effect. Furthermore, as $\Delta$ increases, the initial velocity of the particles decreases, resulting in a longer relaxation time required for the system to return to equilibrium. This indicates that $\Delta$ holds the bulk dynamics of the particles. The phenomenon can be explained by the fact that the smaller the initial velocity of the particles, the longer time it takes for them to reach the edge, and the edge effects exert less influence. In other words, the Wannier-Stark ladder governs the bulk dynamics and influences the relaxation time to equilibrium.
We also computed $\overline{v}$ for different values of $\gamma$. From FIG. 2 (j), it is evident that under $\Delta$ = 0, when $\gamma$ is within an appropriate range (e.g., $\gamma$ = $0.5$, $1.0$, $2.0$ and $4.0$), the evolution of $\overline{v}$ is roughly similar. However, when $\gamma$ is particularly small or particularly large (e.g., $\gamma$ = $0.1$ or $8.0$), the evolution of $\overline{v}$ undergoes significant changes. Unlike variations in $\Delta$, altering $\gamma$ simultaneously changes both $\hat{H}_{\text{eff}}$ and the mean value $\langle\hat{L}_{l}^{\dagger}\hat{L}_{l}\rangle\gamma dt$ of the Poisson random numbers, and the effect of $\gamma$ is non-monotonic.
We investigated the time evolution of additional observables during the relaxation process, with particular emphasis on their behavior at the transition point. In FIG. 3, we display three distinct observables, with each column’s panels corresponding to the same $\Delta$ value. The black dashed line indicates the transition time, $t_{c}/\tau$, which is subsequently validated through finite-size scaling analysis. Panels (a)–(d) illustrate the evolution of $\overline{I}_{AB}$. For various system sizes $L$, we observe that $\overline{I}_{AB}$ initially increases with time, then decreases, eventually reaching zero at $t_{c}/\tau$. This indicates that once the system attains a steady state, there is no correlation between subsystems $A$ and $B$. In panels (e)–(h), we depict the evolution of $\overline{f}_{\text{skin}}$ , which quantifies the probability that the evolving state $|\psi(t)\rangle$ is in the ideal many-body state $|\psi_{\text{skin}}\rangle$. It is evident that before $t_{c}/\tau$, $\overline{f}_{\text{skin}}$ remains at 0, indicating that $|\psi(t)\rangle$ can never be in $|\psi_{\text{skin}}\rangle$. However, at $t_{c}/\tau$ this probability surges sharply, though it does not reach 1. This corresponds to the domain wall state without a distinct boundary, as depicted in FIG. 2 (a) and (b). Panels (i)–(l) depict the evolution of $\overline{\|D\|}_{F}$. This observable, which is similar in nature to $\overline{I}_{AB}$, follows an evolution pattern akin to that of $\overline{S}_{L/2}$. As shown, a crossover occurs at $t_{c}/\tau$, after which the values rapidly stabilize. Taking all of the above observables into account, we find that their critical behavior is also delayed as $\Delta$ increases.
III.3 Criticalities of the dynamical phase transitions
For the free fermion model. the finite size effect is significant, and hence it is not trivial to obtain critical points which separate the log-law phase and the area-law phase in entanglement relaxation dynamics. To investigate this critical properties, we obtained scaling collapses of $\overline{S}_{L/2}$ as a function of rescaled time for $\Delta=0.0$, $0.2$, $0.4$, and $0.6$ under open boundary conditions (OBCs), all of which are shown in the lower zone of FIG. 1. In data collapse analysis, we employed the relation, $\overline{S}_{L/2}(t/\tau,L)L^{-\beta}=F[(t/\tau-t_{c}/\tau)L^{1/\nu}]$, where $F$ is a polynomial function. We identified a set of critical parameters { $t_{c}/\tau$, $\nu$, $\beta$ in Table. 1, which enable data collapse for system of different $L$, shown in FIG. 4.
Table 1: The critical information of dynamical phase transition under different $\Delta$. The initial states in the first three rows correspond to the Néel state, while the initial states in the last row correspond to the anti-Skin-like state (The state where particles accumulate at the opposite boundary).
$\Delta$ $t_{c}/\tau$ $\nu$ $\beta\times 10^{-2}$
$0.0$ $0.79(2)$ $1.23(4)$ $1.57(5)$
$0.2$ $0.89(5)$ $1.20(2)$ $2.69(3)$
$0.4$ $1.10(6)$ $1.19(4)$ $5.20(1)$
$0.6$ $1.45(4)$ $1.12(7)$ $3.12(2)$
$0.0$ $1.46(4)$ $1.23(6)$ $2.89(8)$
Figure 4: Finite-size scaling for $\overline{S}_{L/2}$ with $\overline{S}_{L/2}L^{-\beta}$ versus $(t/\tau-t_{c}/\tau)L^{1/\nu}$. Panels (a)-(d) correspond to $\Delta$ = $0.0$, $0.2$, $0.4$ and $0.6$, respectively. The system sizes $L$ = $32$, $64$, $96$, $128$ and $192$ are represented by a color gradient from light to dark, where lighter shades correspond to smaller $L$ and darker shades to larger $L$. Insets in each panel show the evolution of $\overline{S}_{L/2}$ over time for the corresponding $\Delta$. The black dashed lines in each inset represent the $t_{c}/\tau$ for the respective panel.
It is worth noting that as $\Delta$ increases, the critical time $t_{c}/\tau$ gradually increases as well. However, the critical exponents $\beta$ and $\nu$ remain largely unchanged. Therefore, we conclude that the Wannier-Stark ladder does not alter the intrinsic universality class of this dynamical phase transition. From the perspective of this critical exponent, it suggests that the system belongs to the Ising universality class.
IV Discussion and Conclusion
Here, we have elucidated the mechanism underlying a numerically observed dynamical phase transition in a tilted free fermion chain subject to monitoring-feedback protocols. Our main findings are threefold:
  1. (i)
    Feedback reshapes the system’s entanglement dynamics from a metastable, measurement-induced log-law entangled state to an eventual area-law skin-like steady state.
  2. (ii)
    The resulting transition exhibits finite-size scaling in a dimensionless time variable, $t/\tau$, where $\tau\sim L$ is the characteristic transport time in a 1D open system.
  3. (iii)
    A moderate Wannier-Stark ladder does not affect or alter the universality class of this dynamical phase transition.
Figure 5: Panels (a) and (e) illustrate the schematic diagrams of entanglement structures with and without feedback operations, respectively. Panels (b) and (f) describe the entanglement relaxation process for another initial state. For the initial state, the particles are mainly concentrated on the left side for panel (b) and on the right side for panel (f). The inset in (f) is finite-size scaling for $\overline{S}_{L/2}$ with $\overline{S}_{L/2}L^{-0.029}$ versus $(t/\tau-1.46)L^{1/1.23}$. Panels (c) and (g), we rescale the time and show the evolution of $\overline{S}_{L/2}$. Panels (d) and (h) illustrate the evolution of $\overline{S}_{L/2}$ for different $\gamma$. All panels (b)-(d) and (f)-(h) are in $\Delta=0$.
We begin by analyzing the limit without feedback, $\theta=0$ in Eq. (3). In this regime, the system is effectively a free-fermion setup with a MIET. At measurement rate $\gamma=0.5$, the system resides within the critical MIET phase, where the steady-state entanglement scales logarithmically with subsystem size (log-law). Numerically, this behavior is shown by the blue dashed lines in FIG. 5 (a) and (e), with deeper colors corresponding to larger system sizes. However, in the no-click limit, the dynamics are governed by the non-Hermitian effective Hamiltonian Eq. (4), which drives the system into a skin-like steady state exhibiting area-law entanglement. This purely non-Hermitian behavior is illustrated by the black dashed lines in FIG. 5 (a) and (e).
Separately, introducing the feedback operator Eq. (3) with $\theta=\pi$ balances these two extremes. The feedback acts to erase the noise of quantum jumps, thereby pushing the system toward the area-law entangled, non-Hermitian skin-like steady state. However, the feedback effect is not instantaneous. Initially, the system remains in a metastable log-law entangled regime; only after a characteristic relaxation time does the feedback reshape the evolving state into the final skin-like state. This two-stage process is shown schematically by the solid green lines in FIG. 5 (e) and numerically in FIG 2 (e) (green lines). Early on, the system exhibits log-law entanglement, then crosses over to the area-law entangled steady state. The finite-size crossing in the time domain thus emerges naturally as a feedback-induced dynamical phase transition. Additionally, the lifetime of the metastable log-law regime is determined by how quickly the feedback pumping reshapes the system into the skin-like steady state. If the initial state is already similar to the final steady state, relaxation occurs swiftly, and the metastable window is narrow. Conversely, if the initial state is orthogonal, a longer metastable regime is observed. FIG. 5 (b) illustrates the entanglement dynamics starting from the effective Hamiltonian’s skin-like steady state: no metastable regime appears, as the system is already at equilibrium under the non-Hermitian dynamics. In contrast, FIG. 5 (f) shows the opposite case, where the initial state is a reverse skin state, fully orthogonal to the final steady state. The metastable window is significantly prolonged, and the dynamical transition time increases to $t_{c}/\tau\approx 1.46$, compared to a shorter time in the Néel-state case [FIG 2 (e), green lines]. Since the choice of the initial state does not alter the mechanism between the MIET dynamics and the non-Hermitian effective Hamiltonian dynamics, the critical exponents remain independent of the initial state. Our numerical results are shown in TABLE 1 and the inset of FIG. 5 (f).
To understand the critical behavior, we note that $t/L$ (proposed in Refs. [66, 70, 1, 71]) are dimensionally confusing, as it does not carry units of time. Instead, we identify $t/\tau$ as a more natural scaling variable, where $\tau$ is the time for particles to traverse a 1D open chain of length $L$. As $\tau\sim L$, the ratio $t/\tau$ is dimensionless and suitable for finite-size scaling. When other nonlinear time rescalings are attempted [FIG. 5 (c) and (g)], signatures of the dynamical phase transition disappear, confirming that $t/\tau$ is the appropriate scaling parameter. Moreover, at higher measurement rates $\gamma$ [FIG. 5 (d) and (h)], the metastable state itself becomes area-law, rather than log-law, but we anticipate the same feedback mechanism to induce a crossover between two distinct states in the entanglement evolution. Observing this transition in practice, however, is more challenging because stronger quantum-jump noise obscures the crossover.
It worth to notice that the Wannier-Stark ladder preserves the universality class of this dynamical phase transition. As discussed in Ref. [1], the localization of the effective Hamiltonian only monotonically influences the time it takes for particles to be transported to the boundary. The stronger the localization, the longer the transport time. This suggests that the criticality of this dynamical phase transition is not entirely determined by the dynamics of the effective Hamiltonian.
To sum up, we systematically investigate a dynamical phase transition in monitored free fermion chain subject to feedback control. We study the evolution of the particle density distribution and entanglement entropy under different boundary conditions (open and periodic) and analyze the impact of the Wannier-Stark ladder. Firstly, we identify and characterize the feedback-induced skin effect, where the trajectory-averaged particle density becomes localized on one side of the lattice at late times. Introducing a Wannier-Stark ladder significantly prolongs the residence time in this non-equilibrium state, as the potential gradient restricts particle transport. Under periodic boundary conditions, this gradient creates a pseudo-edge, leading to a similar localization effect even without physical edges. Secondly, by analyzing the trajectory-averaged entanglement entropy, we observe a clear two-stage dynamical process: initially, entanglement rapidly grows into a metastable regime characterized by logarithmic scaling (log-law), and subsequently decays slowly into a final steady state exhibiting area-law entanglement. This transition between log-law and area-law behaviors occurs at a distinct dynamical transition point, confirmed by examining multiple observables, including subsystem mutual information, probability of the evolving state being in the ideal skin state, and subsystem correlation norms. Using finite-size scaling analysis, we precisely identify critical parameters associated with this dynamical phase transition. Importantly, we provide an intuitive and graphical explanation for the mechanism driving this dynamical phase transition: feedback gradually suppresses quantum jump noise, reshaping the system’s steady state from a measurement-induced metastable log-law entangled state into a non-Hermitian skin-like area-law state. Moreover, we discuss how the presence of the Wannier-Stark ladder affects the dynamics: while it significantly delays the relaxation process, it does not alter the universality class of the observed dynamical phase transition. Finally, we explore the fate of this dynamical transition in the presence of a strong Wannier-Stark ladder. We propose that if the gradient is sufficiently large, the system may directly relax into an area-law regime without undergoing a clear dynamical transition.
This work further complements the dynamical phase transition induced by the feedback-induced skin effect previously proposed in Ref. [1]. Additionally, experimentally observing this dynamical phase transition requires a platform capable of precise real-time feedback control, single-particle resolution measurements, and tunable potential gradients, which are likely achievable in ultracold atom optical lattices [94] or ion-trap systems [95, 94, 96]. However, implementing high-precision, rapid feedback loops for dynamically adjusting system parameters, as well as accurately measuring and controlling entanglement entropy and particle density distributions, remains challenging.
@@ Acknowledgements.
We thank Ze-Chuan Liu, Yu-Peng Wang, and Shan-Zhong Li for discussions. H.-Z. Li appreciates the guidance of Ching Hua Lee and Xue-Jia Yu. J.-X. Zhong acknowledges the National Natural Science Foundation of China (Grant No. 12374046 and No. 11874316), the National Basic Research Program of China (Grant No. 2015CB921103), and the Program for Changjiang Scholars and Innovative Research Teams in Universities (Grant No. IRT13093).